# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

### v2

@@ B.2 Deformation to the swindle
Appendix C Homotopy groups of QCA
The $\Omega$-spectrum property reduces the calculation of homotopy groups $\pi_{n}(\mathrm{QCA}_{d})$ of QCA to either finding the connected components $\pi_{0}(\mathrm{QCA}_{d-n})$ or the homotopy groups of a zero-dimensional QCA, $\pi_{n-d}(\mathrm{QCA}_{0})$. In this appendix, we summarize progress in the literature in computing $\pi_{0}(\mathrm{QCA}_{d-n})$ for spin systems (Appendix C.1), and compute all the groups $\pi_{n-d}(\mathrm{QCA}_{0})$ (Appendix C.2). We briefly comment on the classifcation of QCA on lattices other than ${\mathbb{Z}}^{d}$ in Appendix C.3, but leave any technical arguments to Ref. [107].
Unlike the previous two appendices, the results of this appendix are specific for spin systems. The homotopy groups of the space of fermionic QCA will differ from the groups presented here.
C.1 Connected components
In general, the connected components $\pi_{0}(\mathrm{QCA}_{d})$ are not known. However, these groups have been calculated for $d\in\{1,2\}$ (and $d=0$, which is trivial), and the groups are known to be nontrivial in some higher dimensions.
For one-dimensional QCA, the complete classification [that is, $\pi_{0}(\mathrm{QCA}_{1})$] was obtained by Ref. [58], which also explicitly defined an invariant which diagnoses which connected component a given QCA belongs to. All one-dimensional QCA are path connected to a shift—a QCA which acts as a translation on some subset of the degrees of freedom. For example, in a chain of qubits, each QCA is connected to a QCA which translates all qubits by $r$ sites either right or left. For qubit chains, QCA are classified by the group ${\mathbb{Z}}$.
However, we stabilize QCA by allowing ancillae of any Hilbert space dimension to be added to any site by tensor product. Thus, we must assume that the chain also has qutrit (three-state) degrees of freedom, and other higher level qudits. While a 4-level qudit can be decomposed into two qubits, and so does not give new invariants compared to the qubit system, qutrits and other prime qudits all produce invariants. A shift of qutrits cannot be connected to a shift of qubits.
The resulting classifying group is succinctly presented as
$\pi_{0}(\mathrm{QCA}_{1})\cong\log{\mathbb{Q}}^{\times},$ (69)
$\log{\mathbb{Q}}^{\times}\cong\bigoplus_{p\text{ prime}}{\mathbb{Z}}\log p.$ (70)
$\pi_{0}(\mathrm{QCA}_{2})\cong 0.$ (71)
There is some subtlety in this statement. For QCA on a general two-dimensional manifold—that is, the lattice is obtained from a triangulation of a manifold—the classification may be nontrivial. However, all invariants relate to one-dimensional cycles in the manifold, and so belong to a lower dimensional part of the classification. For appropriate choices of topology on $\mathrm{QCA}_{2}$ [107] (where the manifold is ${\mathbb{R}}^{2}$), this allows all two-dimensional QCA on the plane to be connected by continuous paths.
The classification of three dimensional QCA is still open. It is conjectured that $\pi_{0}(\mathrm{QCA}_{3})$ is isomorphic to the Witt group of modular tensor categories [56, 108, 55, 109]—that is, it has the same group structure as stacking two-dimensional topological orders, modulo an equivalence relation which identifies phases which may share a gapped edge. Several nontrivial components of $\pi_{0}(\mathrm{QCA}_{3})$ have been identified, which are believed to form a subgroup [55],
$\pi_{0}(\mathrm{QCA}_{3})\geq{\mathbb{Z}}_{8}\oplus{\mathbb{Z}}_{2}.$ (72)
The elements of this subgroup can, indeed, be identified with two-dimensional topologically ordered phases. In fact, the ALT phases corresponding to these elements exhibit that topological order (specifically, the associated anyonic excitations) on their two dimensional surfaces [48, 56, 55, 109].
C.2 Zero dimensions
$\mathrm{QCA}_{0}(N)\cong PU(N).$ (73)
C.2.1 Unstable invariants
For finite but large $N$, the homotopy groups of $PU(N)$ are all known,
There are even integral formulae for winding number invariants indexing the elements of these groups. Given any continuous map $V:S^{n}\to PU(N)$ (where $S^{n}$ is the $n$-dimensional sphere) it is possible to find a lift $\tilde{V}:S^{n}\setminus\mathrm{pt}\to U(N)$ from the punctured sphere ($\mathrm{pt}$ is the topological space with a single point) to the usual unitary group. That is, if $q:U(N)\to PU(N)$ is the quotient map from the unitary group to the projective unitaries, we have $q\tilde{V}=V|_{S^{n}\setminus\mathrm{pt}}$. This is possible because the punctured sphere is contractible, and so it is possible to give a smooth gauge to $\tilde{V}$ away from the puncture. Defining $A=(\mathrm{d}\tilde{V})\tilde{V}^{\dagger}$, where $\mathrm{d}$ is the exterior derivative on $S^{n}\setminus\mathrm{pt}$, the winding number invariants are given by [31, 130]
$W_{n}[\tilde{V}]=C_{n}\int_{S^{n}\setminus\mathrm{pt}}\mathrm{Tr}\left[A^{\wedge n}\right],$ (75)
$w_{1}[V]=W_{1}[\tilde{V}]\bmod N\in{\mathbb{Z}}_{N}$ (76)
C.2.2 Stable invariants
The complication is that these winding numbers are invariant under the direct sum with an ancilla, not the tensor product. Indeed, denoting $\mathbbm{1}_{M}$ as the identity in $PU(M)$, we have from Eq. (75)
$W_{n}[V\otimes\mathbbm{1}_{M}]=MW_{n}[V].$ (77)
More generally, for $Q\in PU(M)$,
$W_{n}[V\otimes Q]=MW_{n}[V]+NW_{n}[Q].$ (78)
$\tilde{W}_{n}[V]=C_{n}\int_{S^{n}\setminus\mathrm{pt}}\frac{\mathrm{Tr}\left[A^{\wedge n}\right]}{\mathrm{Tr}\left[\mathbbm{1}\right]}.$ (79)
These are no longer integer invariants, but rather rational invariants [modulo 1 for $n=1$]. Then Eq. (78) shows that $\tilde{W}_{n}[V\otimes Q]=\tilde{W}_{n}[V]+\tilde{W}_{n}[Q]$.
C.2.3 Classification
At a more formal level, for all $N,M\in{\mathbb{N}}$ we have injective homomorphisms
$\iota_{N\to NM}:PU(N)$ $\to PU(NM)$ (80)
$V$ $\mapsto V\otimes\mathbbm{1}_{M}$
[…]
with respect to this system. Strictly, we require that the multiplicity $n_{p}$ of every prime $p$ in the factorization $N=2^{n_{2}}3^{n_{3}}\ldots p^{n_{p}}\ldots$ approaches infinity, $n_{p}\to\infty$. The homotopy groups of $\mathrm{QCA}_{0}$ can then be computed from the induced system
$\iota^{*}_{N\to NM}:\pi_{n}(PU(N))$ $\to\pi_{n}(PU(NM))$ (82)
$W_{n}[V]$ $\mapsto MW_{n}[V].$
By rescaling the winding numbers, we find that
[…]
C.2.4 Limited ancillae
We obtained a classification by rationals because we allowed ancillae of any dimension to be appended to the system. Thus, all Hilbert space dimensions can occur in the denominator of $\tilde{W}_{n}$. If, instead, we considered only a limited collection of Hilbert space dimensions for the ancillae, we would obtain a different classification. In fact, this classification will be a useful warm-up for the case of QCA with symmetry.
Say our available ancillae dimensions are given by
=$\mathcal{R}=\{\rho\in{\mathbb{N}}\,:\,\text{ancilla of dimension }\rho\text{ exists}\},$ (88)
C.3 QCA on manifolds
The $\Omega$-spectrum property discussed in Appendix B assumes that QCA are defined on the integer lattice ${\mathbb{Z}}^{d}$. By coarse graining, any locally finite set of points in ${\mathbb{R}}^{d}$ can be deformed to ${\mathbb{Z}}^{d}$, so the QCA is more properly based on the manifold ${\mathbb{R}}^{d}$. This distinction is important, as the classification of QCA based on the $d$-dimensional manifold $M$, $\mathrm{QCA}_{M}$, depends on the homology of $M$ [61] and not just its dimension.
$\mathrm{QCA}_{M}\to\left[M,\mathrm{QCA}_{d}\right]$ (91)
This leads to the hypothesis that, for orientable $M$ [116],
We defer the majority of work to Ref. [107], but we remark here that Eq. (92) reproduces the classification of QCA on two-dimensional orientable manifolds in Ref. [61]. For non-orientable manifolds, one must consider twisted generalized cohomology—that is, $\mathrm{QCA}_{M}$ would be homotopy equivalent to spaces of sections in a $\mathrm{QCA}_{d}$-bundle over $M$ [103, 104].
Appendix D Homotopy groups of symmetric QCA
The QCA formalism for classifying localized phases can be adapted to incorporate symmetry constraints, and thus provide a classification of symmetry-enriched ALT phases (SALT phases), and more generally of symmetry-enriched MBL phases with short-ranged entanglement. In this appendix, we show that the $\Omega$-spectrum property of QCA respects any on-site unitary group action, from which it follows that symmetric QCA also form an $\Omega$-spectrum (Appendix D.1). While the $\Omega$-spectrum property of invertible states is linked to a classification of symmetry-enriched phases by $BG$-parameterized families of states ($BG$ is the classifying space for the symmetry group), this construction fails for QCA (Appendix D.2). Nonetheless, we identify complete invariants for all homotopy groups of symmetric QCA for $d\leq 1$, and give a presentation of the group structure (Appendices D.3D.4).
An important observation is that the classification of symmetric QCA depends on the representation of the symmetry action [93, 26, 114]. This can be formalized within the stabilization procedure of QCA. Each added ancilla carries some representation of the symmetry group $G$. We do not assume that all representations of $G$ are necessarily available. Define a countable set of representations which occur in the ancillae,
:=$\mathcal{R}:=\{\rho\in\mathrm{Rep}(G)\,:\,\text{some ancillae carry }\rho\},$ (93)
where $\mathrm{Rep}(G)$ is the tensor category of finite dimensional unitary representations of $G$. The set $\mathcal{R}$ is a multiplicative submonoid of $\mathrm{Rep}(G)$ under tensor product, as if $\rho$ and $\sigma$ are representations carried by some ancillae, then $\rho\otimes\sigma$ is the representation of the ancilla formed by stacking both ancillae. $\mathcal{R}$ also contains the trivial one-dimensional representation $1$, which corresponds to appending no ancillae. This definition of $\mathcal{R}$ generalizes Eq. (88), for which we can regard each ancilla as a representation of the trivial group. (One can compare $\mathcal{R}$ to the closely related notion of a $G$-universe [131], which is defined similarly, except that a $G$-universe is usually required to be closed under direct sum, rather than tensor product.)
We say that two QCA $V_{1}$ and $V_{2}$ which carry a group action by $G$ are $\mathcal{R}$-equivalent if ancillae from $\mathcal{R}$ can be appended to each of them such that they become equal, and carry the same representation. In the literature, two QCA have been called strongly equivalent if they are connected by a symmetric path modulo $\mathcal{R}$-equivalence [93, 114].
D.1 Symmetric swap
In Appendix B, we argued that the swindle map $S:\mathrm{QCA}_{d-1}\to\Omega\mathrm{QCA}_{d}$ is a homotopy equivalence. Both the swindle map and the deformations which demonstrated this equivalence were defined in terms of a path $\mathrm{swap}_{jk}(t)$ from the identity map to a swap gate between two sites with the same local Hilbert space dimension. In the context of QCA with an on-site $G$-action, we can repeat all the same arguments, provided we can find a path $\mathrm{swap}^{\rho}_{jk}(t)$ which commutes with the representation $\rho$.
where $\mathrm{Sym}(\mathcal{H}_{j}\otimes\mathcal{H}_{k})$ is the $(+1)$-eigenspace of $\mathrm{swap}_{jk}(1)$ and $\mathrm{Sym}(\mathcal{H}_{j}\otimes\mathcal{H}_{k})$ its $(-1)$-eigenspace, we have
$\rho\otimes\rho\cong\rho_{\mathrm{Sym}}\oplus\rho_{\mathrm{Alt}}.$ (96)
Then the path of operators
where $\mathrm{SWAP}^{\rho}_{x,x+1}(t)$ is a product of $\mathrm{swap}^{\rho}_{jk}(t)$ paths, as in Eq. (6). Multiplying these together gives an instance of the swindle map
which respects the group action in the sense that
That is, the group action on the $d$-dimensional loop agrees with the action on the $(d-1)$-dimensional QCA. Further, as the deformation from $\Omega\mathrm{QCA}_{d}$ to the image of $S(v,t)$ in Appendix B proceeded exclusively through interleaving finite depth circuits with swap gates, we can make all the same deformations with $\mathrm{SWAP}^{\rho}_{x,x+1}(t)$. Thus, $S^{\rho}$ gives a $G$-homotopy equivalence—a homotopy equivalence which respects the $G$ action.
This makes $\mathrm{QCA}$ a $G$-spectrum, or more specifically a (naive) $\Omega$-spectrum with $G$-action. Formally, this defines a (Bredon) equivariant cohomology theory. On general grounds, it is known that the fixed point subspaces of the spectrum—QCA which commute with the group action—themselves form an $\Omega$-spectrum [131]. (An equivariant map must map fixed points to fixed points.) We write the fixed point spaces as $\mathrm{QCA}_{d}^{\mathcal{R}}$, where we use $\mathcal{R}$ as the decoration rather than $\rho$ to denote the stabilization by appending ancillae. Also define $\mathrm{QCA}_{d}^{\mathcal{R}}$ for negative $d$ by $\mathrm{QCA}_{d}^{\mathcal{R}}=\Omega^{-d}\mathrm{QCA}_{0}^{\mathcal{R}}$. Then we have
More generally, there is a generalized cohomology theory
which classifies parameterized families of symmetric QCA.
D.2 Symmetric QCA are not classified by $BG$-parameterized QCA
Invertible states are also conjectured to form an $\Omega$-spectrum [96, 95, 102, 103, 104]. Further, it is believed that the classification of $G$-symmetry-enriched invertible states can be deduced just from the cohomology theory $h^{d}_{\mathrm{inv}}(X):=\pi_{0}\left[X,\mathrm{inv}_{d}\right]$ and the abstract group $G$. In particular, denoting $\mathrm{inv}^{G}_{d}$ for the space of fixed points under the $G$-action, it is conjectured that
$\pi_{0}(\mathrm{inv}^{G}_{d})\cong h^{d}_{\mathrm{inv}}(BG),$ (103)
where $BG$ is the classifying space for the group $G$. That is, connected components of symmetric QCA are in bijection with families of not-necessarily-symmetric QCA parameterized by $BG$. This conjecture can has a natural interpretation in which the map in $\pi_{0}\left[BG,\mathrm{inv}_{d}\right]$ encodes the fusion rules of domain walls [96, 95, 113, 102, 103, 104].
The corresponding hypothesis for QCA is, in general, false:
$\pi_{0}(\mathrm{QCA}^{\mathcal{R}}_{d})\not\cong h^{d}(BG).$ (104)
Indeed, the hypothetical classification by $h^{d}(BG)$ is independent of the symmetry representation $\rho$, and as the strong classification of QCA (with restricted representations of ancillae) is known to depend on the representation, it cannot be the case that symmetric QCA are classified by $h^{d}(BG)$. In fact, Ref. [26] computed the classification of $U(1)$-symmetric QCA in one dimension with an arbitrary symmetry representation. In this case, while the classification depends on the symmetric QCA $V$ and the representation $\rho$, the invariant associated to the pair $(V,\rho)$ does not depend on which stabilizing ancillae can be included, $\mathcal{R}$. Thus, some of the subtleties associated to stacking ancillae are absent in this case. Regardless, we already see the $\rho$-independent $h^{1}(BG)$ cannot be a complete classification.
By explicit computation, we show
$\pi_{0}(\mathrm{QCA}^{\mathcal{R}}_{1})\not\geq h^{1}(BU(1))$ (105)
for any collection $\mathcal{R}$ of representations of $U(1)$. Indeed, the left hand side was found in Ref. [26]. We will repeat this classification with greater generality in Appendix D.4, but for now all we need is that the classifying group is free abelian, so no element has finite order.
On the right hand side of Eq. (105), we can compute $h^{1}(BU(1))$ by solving the Atiyah-Hirzebruch spectral sequence (AHSS) [103, 104]. The AHSS is a method for evaluating generalized cohomology theories $h^{d}(X)$ given knowledge of $h^{p}(\mathrm{pt})$ for all $p\leq d+1$ and the cohomology groups of $X$, $H^{p}(X;M)$. In general, it gives that $h^{d}(X)$ is a quotient of a subgroup of
$\bigoplus_{p=0}^{\infty}H^{p}(X;h^{d-p}(\mathrm{pt})).$ (106)
Identifying the right subgroup and quotient is difficult in general, but for $X=BU(1)\simeq{\mathbb{C}}\mathbb{P}^{\infty}$ and $d=1$ it happens to be easy—the spectral sequence stabilizes on the $E_{2}$ page, so the subgroup is the entire direct sum and the quotient is trivial. We have
$h^{1}(BU(1))$ $\cong\bigoplus_{p=0}^{\infty}H^{p}({\mathbb{C}}\mathbb{P}^{\infty};h^{1-p}(\mathrm{pt}))$ (107)
$\cong H^{0}({\mathbb{C}}\mathbb{P}^{\infty};\log{\mathbb{Q}}_{+})\oplus 0\oplus H^{2}({\mathbb{C}}\mathbb{P}^{\infty};{\mathbb{Q}}/{\mathbb{Z}})$
$\qquad\qquad\oplus 0\oplus H^{4}({\mathbb{C}}\mathbb{P}^{\infty};{\mathbb{Q}})\oplus 0\cdots$ (108)
$\cong\log{\mathbb{Q}}^{\times}\oplus{\mathbb{Q}}/{\mathbb{Z}}\oplus{\mathbb{Q}}\oplus{\mathbb{Q}}\oplus\cdots.$ (109)
We used that $BU(1)={\mathbb{C}}\mathbb{P}^{\infty}$ and $H^{n}({\mathbb{C}}\mathbb{P}^{\infty};M)=M$ for even $n$ and is trivial otherwise. In particular, we have $h^{1}(BU(1))\geq{\mathbb{Q}}/{\mathbb{Z}}$, which is not a free group—all its elements have finite order. Thus, $\pi_{0}(\mathrm{QCA}^{\mathcal{R}}_{1})$ cannot contain $h^{1}(BU(1))$.
In the case of invertible states, the classification by $h^{d}_{\mathrm{inv}}(BG)$ is justified through a decorated domain wall construction [103, 104]. We speculate that this fails for QCA because, while QCA form a spectrum, that spectrum is non-connective—it has nontrivial homotopy groups $h^{d}(\mathrm{pt})$ in arbitrarily negative degree $d<0$. Thus, the putative domain wall structure encoded in $h^{d}(BG)$ includes algebraic fusion rules for domain walls which exist in negative dimension, which are not physical. This may also be why single-particle symmetric phases are not classified by $h^{d}_{K}(BG)$ even though the phases without symmetry are classified by the $K$-theory spectrum $h^{d}_{K}$ (which has nontrivial homotopy groups in all even degrees) [100].
It seems plausible that all symmetric QCA can be classified through a decorated domain wall construction. Certainly, all examples in the literature have this structure [91, 26, 93]. However, when correctly implemented, this approach must treat the zero-dimensional defects which terminate such constructions [113] with greater care than $h^{d}(BG)$.
D.3 Zero dimensions with symmetry
As in the case without symmetry, the homotopy groups $\pi_{n}(\mathrm{QCA}^{\mathcal{R}}_{d})$ with $d\leq n$ are determined by $\pi_{n-d}(\mathrm{QCA}^{\mathcal{R}}_{0})$. Thus, we make significant progress on characterizing symmetric QCA in higher dimensions through an in-depth analysis of zero-dimensional QCA with symmetry. Our strategy is the same as the case without symmetry: we identify unstable invariants (at fixed Hilbert space dimension) and subsequently find a stable presentation of these invariants which does not change upon adding ancillae. Finally, we characterize the abstract classifying group.
D.3.1 Unstable invariants
@@
$\tilde{V}=\bigoplus_{r\text{ irrep of }G}\mathbbm{1}_{d_{r}}\otimes\tilde{V}_{r}.$ (112)
@@
which consists of finite formal sums of winding numbers with coefficients in $\mathrm{Rep}(G)$, regarded as a rig (a ring without negatives). That is, addition in $\mathrm{Rep}(G)$ is direct sum, multiplication is tensor product (which distributes over direct sums) and elements are regarded up to unitary equivalence of representations, so that addition and multiplication are commutative. We do not append the formal negatives used to define the representation ring $R_{G}$. However, as $\pi_{n}(U(N))\in\{0,{\mathbb{Z}}\}$ for sufficiently large $N$, the module $\mathrm{Rep}(G)[\pi_{n}(U(N))]$ is isomorphic to either 00 or $R_{G}$.
@@
QCA are properly projective unitaries, rather than unitaries. $W^{\rho}_{n}[\tilde{V}]$ continues to function as a complete invariant for $n\geq 2$, as in the case without symmetry. We may thus denote them $W^{\rho}_{n}[V]$, where $V:S^{n}\to PU(N)$ is a family of symmetric projective unitaries which lie in the connected component of the identity. However, there are symmetric projective unitaries $V$ which cannot be expressed as symmetric proper unitaries (requiring a larger group of $n=0$ invariants), and there are loops $V(\theta)$ of symmetric proper unitaries which become equivalent when regarded as symmetric projective unitaries (requiring a smaller group of $n=1$ invariants).
We first recall that connected components of symmetric projective unitaries [$n=0$] are labelled by one-dimensional representations of $G$. Given a symmetric projective unitary $V$, we can arbitrarily choose a proper unitary $\tilde{V}$ in the preimage of $V$ under the quotient by global phases. The fact that $V$ is symmetric only implies that $\tilde{V}$ is symmetric up to a global phase,
$\rho(g)^{\dagger}\tilde{V}\rho(g)=\chi[V](g)\tilde{V}.$ (116)
Note that $\chi[V]$ is the same when $\tilde{V}$ is multiplied by a global phase, justifying the notation $\chi[V]$ rather than $\chi[\tilde{V}]$. Using that $\rho$ is a representation, we see that $\chi[V]:G\to U(1)$ is a one-dimensional representation of $G$. From Eq. (110), we see that $\tilde{V}$ must decompose as
$\tilde{V}=\sum_{r\text{ irrep of }G}\sigma_{r\to\chi[V]r}\otimes\tilde{V}_{r},$ (117)
where $\sigma_{r\to\chi[V]r}$ is a matrix with a single nonzero element equal to 1, such that $\tilde{V}$ maps the irrep $r$ block of $\rho$ to the irrep $\chi[V]r$ block. Any path of projective unitaries $V(t)$ lifts to a path of proper unitaries $\tilde{V}(t)$ (as the unit interval is contractible), and the space of permutations $\sigma_{r\to\chi[V]r}$ is discrete, so the permutation label $\chi[V]$ is an invariant for path components of projective unitaries. Any unitaries with the same $\chi[V]$ in the decomposition Eq. (117) can be connected by a path, simply by connecting each $\tilde{V}_{r}$ by a path. Thus, $\chi[V]$ is a complete invariant.
@@
As an aside related to the discussion in Appendix D.2, we note that
$X_{G}^{T}\cong H^{1}(G;{\mathbb{Q}}/{\mathbb{Z}})\cong H^{1}(G;\pi_{1}(\mathrm{QCA}_{0}))$ (119)
for the vast majority of groups $G$ relevant to physics. For instance, a sufficient condition is that the component group $G/G^{0}$ (where $G^{0}$ is the identity component of the topological group $G$) is finitely generated.
@@
$W^{\rho}_{1}[e^{im\theta}\tilde{V}]$ $=\sum_{r\text{ irrep of }G}r\cdot W_{1}[e^{im\theta}\tilde{V}_{r}]$ (120)

### v3

@@ B.2 Deformation to the swindle
Appendix C Higher homotopy groups of symmetric QCA
The symmetric implementation of the swap gate constructed in Sec. IV.1 allows all the arguments demonstrating the $\Omega$-spectrum property to be carried out identically for symmetric QCA. As in the case without symmetry, the homotopy groups $\pi_{n}(\mathrm{QCA}^{\mathcal{R}}_{d})$ with $d\leq n$ are determined by $\pi_{n-d}(\mathrm{QCA}^{\mathcal{R}}_{0})$. Thus, we make significant progress on characterizing symmetric QCA in higher dimensions through an in-depth analysis of zero-dimensional QCA with symmetry.
We will assume that the symmetry acts linearly on the local degrees of freedom. Indeed, as we are concerned with the stable classification, we can always consider stacking with an ancilla representation such that the symmetry becomes linear.
In Appendix C.1 we compute all the homotopy groups of symmetric QCA in zero dimensions. We verify a functorial property of the classification in Appendix C.2, and apply the classification to the QP pump of Sec. V.3 in Appendix C.3.
C.1 Zero dimensions with symmetry
Our strategy to classify higher homotopy groups of QCA with symmetry is the same as the case without symmetry: we identify unstable invariants (at fixed Hilbert space dimension) and subsequently find a stable presentation of these invariants which does not change upon adding ancillae. Finally, we characterize the abstract classifying group.
C.1.1 Unstable invariants
@@
[…]
@@
which consists of finite formal sums of winding numbers with coefficients in $\mathrm{Rep}(G)$, regarded as a rig (a ring without negatives). That is, addition in $\mathrm{Rep}(G)$ is direct sum, multiplication is tensor product (which distributes over direct sums) and elements are regarded up to unitary equivalence of representations, so that addition and multiplication are commutative. We do not append the formal negatives used to define the representation ring $R_{G}$. However, as $\pi_{n}(U(N))\in\{0,{\mathbb{Z}}\}$ for sufficiently large $N$, the module $\mathrm{Rep}(G)[\pi_{n}(U(N))]$ is isomorphic to either 0 or $R_{G}$.
@@
QCA are properly projective unitaries, rather than unitaries. $W^{\rho}_{n}[\tilde{V}$] continues to function as a complete invariant for $n\geq 2$, as in the case without symmetry. We may thus denote them $W^{\rho}_{n}[V]$, where $V:S^{n}\to PU(N)$ is a family of symmetric projective unitaries. However, there are symmetric projective unitaries $V$ which cannot be expressed as symmetric proper unitaries (requiring a larger group of $n=0$ invariants), and there are loops $V(\theta)$ of symmetric proper unitaries which become equivalent when regarded as symmetric projective unitaries (requiring a smaller group of $n=1$ invariants).
We first recall that connected components of symmetric projective unitaries ($n=0$) are labelled by one-dimensional representations of $G$. Given a symmetric projective unitary $V$, we can arbitrarily choose a proper unitary $\tilde{V}$ in the preimage of $V$ under the quotient by global phases. The fact that $V$ is symmetric only implies that $\tilde{V}$ is symmetric up to a global phase,
$\rho(g)^{\dagger}\tilde{V}\rho(g)=\chi[V](g)\tilde{V}$, (174)
where we assume that $\chi[V](g)$ is continuous as a function of $g$. Note that $\chi[V]$ is the same when $\tilde{V}$ is multiplied by a global phase, justifying the notation $\chi[V]$ rather than $\chi[\tilde{V}$]. Using that $\rho$ is a representation, we see that $\chi[V]:G\to U(1)$ is a one-dimensional representation of $G$. From Eq. (168), we see that $\tilde{V}$ must decompose as
[…]
where $\sigma_{r\to\chi[V]r}$ is a matrix with a single nonzero element equal to 1, such that $\tilde{V}$ maps the irrep $r$ block of $\rho$ to the irrep $\chi[V]r$ block. Any path of projective unitaries $V(t)$ lifts to a path of proper unitaries $\tilde{V}(t)$ [as the quotient $q:U(N)\to PU(N)$ is a fibration], and the space of permutations $\sigma_{r\to\chi[V]r}$ is discrete, so the permutation label $\chi[V]$ is an invariant for path components of projective unitaries. Any unitaries with the same $\chi[V]$ in the decomposition Eq. (175) can be connected by a path, simply by connecting each $\tilde{V}_{r}$ by a path. Thus, $\chi[V]$ is a complete invariant.
@@
[…]